OpenStudy (praxer):

$$\frac{xdx+ydy}{xdy-ydx}=\sqrt{\frac{a^2-x^2-y^2}{x^2+y^2}}$$ solution: $$sin^{-1}(\frac{\sqrt{x^2+y^2}}{a})=tan^{-1}(\frac{y}{x})+c$$

OpenStudy (praxer):

@sidsiddhartha please help me with this question.

OpenStudy (sidsiddhartha):

hey @praxer how are u? :)

OpenStudy (praxer):

I am fine but this question is like eating me up. Please help me out. :)

OpenStudy (sidsiddhartha):

os is lagging too much here, wait a bit

OpenStudy (sidsiddhartha):

\[x=rcos \theta\\y=r \sin \theta\]

OpenStudy (sidsiddhartha):

again\[x^2+y^2=r^2\\differentiating\\x+y.\frac{ dy }{ dx}=r \frac{ dr }{ dx }\\x.dx+y.dy=r.dr\] ok?

OpenStudy (praxer):

got it :) than for the denominator substitute y/x = v ?????

OpenStudy (sidsiddhartha):

yes again u can write \[\tan \theta=y/x\\so\\ \frac{ x.dy-ydx }{ x^2 }=\sec^2 \theta.d \theta\\x.dy-y.dx=r^2.d \theta\]

OpenStudy (sidsiddhartha):

now just substitute them-- \[\frac{ r.dr }{ r^2.d \theta }=\sqrt{\frac{ 1-r^2 }{ r^2 }}\]

OpenStudy (sidsiddhartha):

\[\frac{ dr }{ \sqrt{1-r^2} }=d \theta\\sin^{-1}r=\theta+c\]

OpenStudy (sidsiddhartha):

\[\sin^{-1} \sqrt{x^2+y^2}=\tan^{-1}y/x+c\]

OpenStudy (sidsiddhartha):

got it?

OpenStudy (praxer):

$$\color{red}{Thank \ You}$$

OpenStudy (sidsiddhartha):

yw!

OpenStudy (sidsiddhartha):

gotta go now :) good night !!

OpenStudy (praxer):

Good night :) Lord be with you :)